Given P = 24 – 0.1Q, where Q = q1+q2 and q1 = q2, TC1 = 0.1q1 2 , TC2 = 0.05q2 2 , a) Determine the output and price of low-costfirm b) Calculate the profit of the low-cost firm c) What isthe profit maximizing price level the high firm would like to charge but that doesn’t realize in the market? d) Compare the profits of the price taker at its own profit maximizing output and low-cost firm’s output e) Show the results a to d graphically 


1
Expert's answer
2022-03-16T18:34:51-0400

profit and cost is maximised when MR=MC

p=24-0.1q

TR=PQ

=(24-.1Q)q

TR=24Q-0.1"q^2"

MR="\frac{d(TR)}{dQ}" =24-0.2q


TC=TC1+TC2

TC=0.1q1"^2" +0.05q2"^2"

MC1="\frac{d(TC1)}{dQ1}" =0.2q1

MC2="\frac{d(T2)}{d2}" =0.1q2


mr1=mc1

24-0.2q1=0.2q1

24=0.4q1

q1=60


mr2=mc2

24-0.2q2=0.1q2

24=0.3q2

q2=80

remember Q=q1+q2

Q=60+80

Q=140

P=24-0.1Q

P=24-(0.1"\times140)"

p=$10


b) profit=TR-TC

TR=QP

=140"\times" 10

=$1400


Tc1=0.1"\times"60"^2"

=0.1"\times" 3600

=$360


tc2=0.05"\times" 80"^2"

=0.05"\times" 6400

=$320

TC=360+320

=680

TC-TR

=1400-680

Profit=$720


c)

price=profit divide by quantity

"\frac{720}{140}" =$5.14